OpenStudy (anonymous):

Solve the following system. 4x^2 + 9y^2 = 72 x - y^2 = -1

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow -y^2 = -1 - x }\) \(\Large \color{MidnightBlue}{\Rightarrow y^2 = 1 + x }\) Solve it by substitution.

OpenStudy (apoorvk):

so, x - y^2 = -1 or, y^2 = x+1 substitute this value of y^2 in the the first equation, and solve the quadratic in x hence obtained.